Given that for an AP seventh term=43 and second term=91,find the valves of a,d and n.
Answer by richwmiller(17219)

an = a + (n - 1)d
43 = a + 6d
91 = a + d
-48=5d
-48/5=d
91 = a -48/5
a = 503/5, d = -48/5
n is whatever you want it to be.